(a+b) is proportional to c. And (b+c) is proportional to a. Prove that (a+c) is proportional to b.

(a+b)  is proportional to c.
And
(b+c) is proportional to a.
Prove that
(a+c) is proportional to b.

Grade:10

1 Answers

Saumya
4 years ago
a+b = kc
b+c = ka
Adding both,
(a+c)+2b = k(a+c)
2b = (a+c)(k-1)
a+c = 2b/(k-1)
So, a+c is proportional to b.